OpenStudy (mackenzie2013):

Simplify -6 squared ÷ 12 - 2(-7). -17 35 11 -35

OpenStudy (anonymous):

none of these are right, -6^2 is 36, 36/12 =3 and then 2(-7) =-14, 3-14= -11

OpenStudy (anonymous):

36/12 + [-2*-7] =3+14 = 17? where is the correct answer?

OpenStudy (mackenzie2013):

so the -2*-7 becomes a positive .. ahh gotha thank you

OpenStudy (mackenzie2013):

wait no

OpenStudy (mackenzie2013):

none of these awnsers are correct its either -17 or 11

OpenStudy (anonymous):

what do you mean if I do what spidy did its 3-(-14) = 3+14 = 17

OpenStudy (anonymous):

are those your only options are confident.

OpenStudy (mackenzie2013):

but thats not a option

OpenStudy (mackenzie2013):

these are the only options

OpenStudy (anonymous):

 perhaps there are some more symbols in the problem. could you use equal editor to write the problem. exacly